1/1

Friedrich Schiller,  Friedrich Schiller,
La distiko  Das Distichon

translated by Manfred Retzlaff

En heksametro sin levas la flua kolon’ de l’ fontano, En pentametro jen falas ĝi laŭ melodi’.

Im Hexameter steigt des Springquells flüssige Säule, Im Pentameter drauf fällt sie melodisch herab.

Translation of the German poem “Das Distichon” by Friedrich Schiller (Frederiko Ŝilero, ∗1759 – †1805) into Esperanto by Manfred Retzlaff (∗1938-11-04) on 2013-04-28.

Arg-891-1774 (2013-04-28 16:30:43)

Author of this German poem is Friedrich Schiller (Frederiko Ŝilero, ∗1759 – †1805).

Arg-891-1773 (2013-04-28 15:49:29)

http://www.poezio.net/version?poem-id=891&version-id=,1774,1773